Larry borrows 18300 dollars from Moe at an effective rate of 9.5 percent, and agrees to make 12 equal annual payments (the first a year from now) to repay the loan. Immediately after Larry makes the seventh payment, Moe sells the loan to Curly. If Moe's total yield rate is 7 percent effective, how much does Curly pay for the loan?
Answer #1

Annual payment=loan*rate/(1-1/(1+rate)^number of payments)=18300*9.5%/(1-1/(1+9.5%)^12)=2620.335171

Curly will pay=2620.335171/7%*(1-1/(1+7%)^5)=10743.89155